# Modelling Gaia photometry signals of dark halos

### v1

@@ 4 Dark halo light curve (Black hole with a dark halo)
$u=x-\frac{1}{x}\left\{\begin{tabular}[]{cc}$1$&$,x\geq x_{R}$\\ $1-q^{-1}\left[1-(x/x_{R})^{2}\right]^{3/2}$&$,x<x_{R}$\end{tabular},\right.$ (17)
@@
5 Detection probability with the Gaia sampling
@@ 6 Summary and discussion
At this point, we should emphasise an important difference between isolated DHs sought-for in surveys, thus far (Alcock et al., 2000; Tisserand et al., 2007; Wyrzykowski et al., 2011; Mróz et al., 2024), and the DHs that are parts of binary systems expected in Gaia’s data. As pointed out by Croon et al. (2020a), the lenses where one can have a detectable signal in the former case, are those with a size comparable to $\theta_{E}$ (i.e., $x_{R}\approx 1$, see Figures 4 and 6). In the latter case though, one can have access to less compact systems ($x_{R}\gg 1$). Gaia’s astrometric observations can measure the masses of the binary members, as well as their orbits (size and inclination) within some observational precision. Knowing the mass of the lens and the geometry of the lens-light source system can allow for measurements of weaker signals, i.e., the difference between the expected brightening for a point mass lens from the observed brightening due to a DH, $|m-m_{pm}|$, for a given mass (see Figure 8).
Figure 8: Plots of the absolute difference between the expected magnitude due to lensing by a point mass lens, $m_{pm}$, and the magnitude $m$ due to lensing by a DH (dashed) and a DH+BH (solid) system of the same mass.
One can see that in the case of a DH-only lens, there is a difference from the expected brightness from a point-mass-lens for $u\rightarrow 0$, which is due to the diffuse DH mass providing a lot less magnification, while for a DH+BH lens, there is a difference because the BH at the centre of the DH, lenses the light source as a point-mass-lens but with less mass than the total mass of the DH+BH system as estimated by the orbital dynamics inferred by the astrometry. One could say that there will be a mass discrepancy between the astrometry and the weak lensing in the latter case. Thus Gaia systems will give access to DH systems not available to previous surveys.
This, however, is the first step, relying on simplifying assumptions, such as a DH of constant density and an ideal orientation of the binary’s plane with respect to the observer. When confronted with observations, these assumptions must be challenged. For an analysis with additional mass profiles for the DH we refer the reader to Romao & Croon (2024).
@@
In conclusion, we need to emphasise an important aspect of binary systems with DHs. The observation of specific, astrometrically-resolved binaries is fundamentally different to previous probes which focused on searching for serendipitous, one-time events. The Gaia DH systems will be available to observe repeatedly (in addition, we will be able to anticipate the occurrence of these lensing events with relative accuracy), thus allowing for higher precision measurements of the structure of the DH.
@@ Acknowledgments
We are indebted to Benoit Famaey for invaluable suggestions, and thank Bradley Kavanagh for discussions. We also thank Lukasz Wyrzykowski for very useful discussions and his feedback. The publication of the article in OA mode was financially supported by HEAL-Link. No new data were generated or analysed in support of this research.
Appendix A BH and DH system

### v2

@@ IV Dark halo light curve (Black hole with a dark halo)
[…]
@@
V Detection probability with the Gaia sampling
@@ VI Summary and discussion
Figure 8: Plots of the absolute difference between the expected magnitude due to lensing by a point mass lens, $m_{pm}$, and the magnitude $m$ due to lensing by a DH (dashed) and a DH+BH (solid) system of the same mass. The spikes in the curves are due to the fact that the $y$-axis is logarithmic and there are some values of $\beta/\theta_{E}$ for which $|m-m_{pm}|=0$. The magnification $\mu$ for the DH or the BH+DH far from the lens is always a little smaller than that of the point mass. Then, near the surface of the DH or the DH+BH, $d\tilde{\alpha}/dx$ becomes a little less steep than what it is for the point mass lens (see Figures 2 and 5), raising $dx/du$ closer to the value of 1 (see eqs. 1011) thus making $\mu$ larger from that of a point mass, which introduces the first outermost crossing near $\beta/\theta_{E}=x_{R}$. Then further in, the point mass $\mu$ becomes larger again, as explained in the text, therefore causing the second crossing.
@@
At this point, we should emphasise an important difference between isolated DHs sought-for in surveys, thus far (Alcock et al., 2000; Tisserand et al., 2007; Wyrzykowski et al., 2011; Mróz et al., 2024), and the DHs that are parts of binary systems expected in Gaia’s data. As pointed out by Croon et al. (2020a), the lenses where one can have a detectable signal in the former case, are those with a size comparable to $\theta_{E}$ (i.e., $x_{R}\approx 1$, see Figures 4 and 6). In the latter case though, one can have access to less compact systems ($x_{R}\gg 1$). Gaia’s astrometric observations can measure the masses of the binary members, as well as their orbits (size and inclination) within some observational precision. Knowing the mass of the lens and the geometry of the lens-light source system can allow for measurements of weaker signals, i.e., the difference between the expected brightening for a point mass lens from the observed brightening due to a DH, $|m-m_{pm}|$, for a given mass (see Figure 8). One can see that in the case of a DH-only lens, there is a difference from the expected brightness from a point-mass-lens for $u\rightarrow 0$, which is due to the diffuse DH mass providing a lot less magnification, while for a DH+BH lens, there is a difference because the BH at the centre of the DH, lenses the light source as a point-mass-lens but with less mass than the total mass of the DH+BH system as estimated by the orbital dynamics inferred by the astrometry. One could say that there will be a mass discrepancy between the astrometry and the weak lensing in the latter case. Thus Gaia systems will give access to DH systems not available to previous surveys.
This, however, is the first step, relying on simplifying assumptions, such as a DH of constant density, an ideal orientation of the binary’s plane with respect to the observer and a point size source. When confronted with observations, these assumptions must be challenged. For example, the finite size of the light source, i.e., the companion star, is something that needs to be taken into account in order to have a realistic model for the light curve from these binaries Witt and Mao (1994); Matsunaga and Yamamoto (2006). The finite size of the light source is expected to smooth out the sharp features of the magnification profiles for the DH systems, as it does for the point mass. However, depending on the situation, either the total magnification or the shape of the light curve, or both, are expected to be different, in particular due to the features that exist near the outer parts of the DH distributions. These features are, for example, expected to modify the light curve during the start and the end of a transit, while the extended nature of the DH will affect the light curves in a way distinctive from that of a point mass even in the case when the center of the lens is well outside the apparent disk of the star. All these effects will be explored in detail in a forthcoming investigation. Furthermore, for an analysis with additional mass profiles for the DH we refer the reader to Romao and Croon (2024).
@@
In conclusion, we need to emphasise an important aspect of binary systems with DHs and stelar companions. The observation of specific, astrometrically-resolved binaries is fundamentally different to previous probes which focused on searching for serendipitous, one-time events. The Gaia DH binary systems will be available to observe repeatedly as the star orbits the DH lens (in addition, we will be able to anticipate the occurrence of these lensing events with relative accuracy), thus allowing for higher precision measurements of the structure of the DH.
@@ Acknowledgments
We are indebted to Benoit Famaey for invaluable suggestions, and thank Bradley Kavanagh for discussions. We also thank Lukasz Wyrzykowski for very useful discussions and his feedback. We also thank the referee for their useful feedback.
Appendix: BH and DH system